# An algorithmic version of the Hajnal–Szemerédi theorem

### v2

@@ 6.5. Auxiliary lemmas
There is an algorithm with running time $n^{O(1)}2^{O(c^{{r}^{r+2}})}$ which either outputs an $I^{*}$-admissible $K_{r}$-tiling $K^{1}$ or returns $(A_{1}^{1},A_{2}^{1},\ldots,A_{s}^{1},B^{1})$ as a certificate showing that $G$ has no $I^{*}$-admissible $K_{r}$-tiling (and thus also no $K_{r}$-factor).
@@
Suppose that $I\neq[s+1]$ and Lemma 6.12 outputs an admissible $K^{1}$. Then we can always find an $[s+1]$-admissible $K_{r}$-tiling $K$ of size at most $r^{2}\mu^{2/3}n$ in time $O(n^{r})$ such that $K^{1}\subseteq K$ and $V(G)\setminus V(K)$ has a partition $A_{1}^{2},\ldots,A_{s}^{2},B^{2}$ with $|A_{1}^{2}|=\cdots=|A_{s}^{2}|=\frac{|B^{2}|}{r-s}\geq n/r-r^{2}\mu^{2/3}n$, $A_{i}^{2}\subseteq A_{i}^{1}$ for each $i\in[s]$ and $B^{2}\subseteq B^{1}$. In particular, if $|B^{1}\setminus V(K)|-(r-s)(n/r-|K^{1}|)>0$, then $K$ contains a copy of $K_{r}$ that has exactly $r-s+1$ vertices of $B^{1}$.
@@
We search $K_{r}$-tilings $F$ of size at most
$(3t_{I})^{r}\leq(3r(rc/0.33)^{r^{r+1}})^{r}\leq((r^{2}c/0.11)^{r^{r+1}})^{r}=(r^{2}c/0.11)^{r^{r+2}}$
(by (I3)) in $G$ such that $\textbf{s}(F)|_{i}=t_{i}$ for each $i\in I^{*}$. By Lemma 1.7, if there is no such $K_{r}$-tiling, then there is no $K_{r}$-factor in $G$. To find such a $K_{r}$-tiling or a certificate showing that none exists, we do the following two steps:
  • Step 1. Search $K_{r}$-tilings of size at most $(3t_{I})^{r}\leq(r^{2}c/0.11)^{r^{r+2}}$ in $G$.
@@
Step 1. Recall from Definition 5.1 that if we consider the partition $(A_{1}^{1},\dots,A_{s}^{1},B^{1})$, then there are $y:={r+s\choose s}$ types of copies of $K_{r}$ in $G$. We fix an arbitrary ordering of these $y$ types. Let $x_{i}$ be the number of copies of $i$th-type $K_{r}$ in a $K_{r}$-tiling of size at most $(3t_{I})^{r}$. Thus if we only concern about the types of $K_{r}$’s (that is, identify copies of $K_{r}$ of same type), then the number of distinct $K_{r}$-tilings of size at most $(3t_{I})^{r}$ is equal to the number of solutions of $\sum_{i=1}^{y}x_{i}\leq(3t_{I})^{r}$, i.e. the number of solutions of $\sum_{i=1}^{y}x_{i}+x_{y+1}=(3t_{I})^{r}$ with $x_{1},\dots,x_{y+1}\in\{0\}\cup\mathbb{N}$, which is ($(3t_{I})^{r}+y\choose y$. For each solution $(x_{1},x_{2},\ldots,x_{y})$ of $\sum_{i=1}^{y}x_{i}\leq(3t_{I})^{r}$, we can decide whether $G$ has a $K_{r}$-tiling of size $C:=\sum_{i=1}^{y}x_{i}$ in time $O(n^{r+1}2^{O(rC)})=O(n^{r+1}2^{O(c^{{r}^{r+2}})})$ by Lemma 5.2. Since there are (${(3t_{I})^{r}+y\choose y}=O(c^{{2r}^{2r+1}})$ solutions, the running time of this step is $O(c^{{2r}^{2r+1}})\cdot O(n^{r+1}2^{O(c^{{r}^{r+2}})})=n^{O(1)}2^{O(c^{{r}^{r+2}})}$ as $c=o(n^{1/{r^{r+3}}})$.
Step 2. Note that checking the slack of a $K_{r}$-tiling takes $O(n)$ time. Since Step 1 returns at most ($(3t_{I})^{r}+y\choose y$ $K_{r}$-tilings, the running time of this step is $O(c^{{2r}^{2r+1}})\cdot O(n)=n^{O(1)}$. If there is a $K_{r}$-tiling $F$ obtained in Step 1 satisfying $\textbf{s}(F)|_{i}=t_{i}$ for each $i\in I^{*}$, then we take it as $K^{1}$. Otherwise, by Lemma 1.7, there is no $K_{r}$-factor in $G$, then we halt and output the partition $(A_{1}^{1},\ldots,A_{s}^{1},B^{1})$ of $G$ as a certificate.
Note that the total running time of the above two steps is $n^{O(1)}2^{O(c^{{r}^{r+2}})}$. ∎
@@
Proof of Lemma 6.15.
It suffices to prove that there is an edge even after deleting any set $U\subseteq A_{i}^{1}$ of size $2(xt_{i}-1)$. Since $|A_{i}^{1}|=\frac{n}{r}+t_{i}$, we have $\delta(G[A_{i}^{1}])\geq t_{i}-c$. On the one hand, $e(G[A_{i}^{1}])\geq\frac{1}{2}|A_{i}|\delta(G[A_{i}^{1}])\geq\frac{1}{2}(\frac{n}{r}+t_{i})(t_{i}-c)\geq\frac{n}{2r}(t_{i}-c)$. On the other hand, $2(xt_{i}-1)\Delta(G[A_{i}^{1}])\leq 2(xt_{i}-1)(3\mu^{1/3}n+2c)\leq 2xt_{i}\cdot 6\mu^{1/3}n$, as $c=o(n^{1/r^{r+3}})$. Hence there is an edge in $G[A_{i}^{1}\setminus U]$ as long as $\frac{n}{2r}(t_{i}-c)>2xt_{i}\cdot 6\mu^{1/3}n$, that is, $x<1/(2400r\mu^{1/3})$. ∎
@@
Let ${A_{i}^{1}}^{\prime}=A_{i}^{1}\setminus V(K^{1})$ and ${B^{1}}^{\prime}=B^{1}\setminus V(K^{1})$. Let $(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{s+1}^{\prime})$ be the slack of partition $({A_{1}^{1}}^{\prime},\ldots,{A_{s}^{1}}^{\prime},{B^{1}}^{\prime})$. Then we have $t_{i}^{\prime}=0$ for each $i\in I^{*}$ and $\sum_{i\in[s+1]\setminus I^{*}}t_{i}^{\prime}=0$. Note that $|K^{1}|\leq(3t_{I})^{r}$.
For each $i\in[s]\setminus I^{*}$, we have $t_{i}^{\prime}\leq t_{i}+|K^{1}|\leq t_{i}+(3t_{I})^{r}$ (a copy of $K_{r}$ in $K^{1}$ can only make the slack of $A_{i}^{1}$ increase by at most 1). Recall that $|t_{i}|\geq\max\{(3t_{I})^{r},c/0.99\}$ by (I2). Hence, if $t_{i}\geq\max\{(3t_{I})^{r},c/0.99\}$, then $t_{i}^{\prime}\leq 2t_{i}$; if $t_{i}\leq\min\{-(3t_{I})^{r},-c/0.99\}$, then $t_{i}^{\prime}\leq 0$. By Lemma 6.15, if $t_{i}\geq\max\{(3t_{I})^{r},c/0.99\}$, then $A_{i}^{1}$ has a matching of size at least $(r+2)t_{i}$ and so $A_{i}^{1^{\prime}}$ has a matching of size at least $2t_{i}$. Hence if $t_{i}^{\prime}>0$, then we can find a matching $M_{i}$ of size $t_{i}^{\prime}$ in time $O(n^{4})$ by the Edmonds algorithm.
If $s+1\notin I^{*}$, then we have $t_{s+1}^{\prime}\leq t_{s+1}+(r-s)|K^{1}|$. Hence, if $t_{s+1}\geq(3t_{I})^{r}$, then $t_{s+1}^{\prime}\leq(r-s+1)t_{s+1}<s(r-s+1)\mu^{2/3}n$, as $t_{s+1}<s\mu^{2/3}n$; if $t_{s+1}\leq-(3t_{I})^{r}$, then $t_{s+1}^{\prime}\leq(r-s-1)(3t_{I})^{r}\leq(r-s-1)(r^{2}c/0.11)^{r^{r+1}}$, as $(3t_{I})^{r}\leq(r^{2}c/0.11)^{r^{r+1}}$. By Lemma 6.9, there are at least $\mu^{7/12}n^{r-s+1}$ copies of $K_{r-s+1}$ in $G[B_{g}^{1}]$, and so $G[B_{g}^{1^{\prime}}]=G[B_{g}^{1}\setminus V(K^{1})]$ contains at least $\mu^{7/12}n^{r-s+1}-r|K^{1}|n^{r-s}\geq(1/2)\mu^{7/12}n^{r-s+1}$ copies of $K_{r-s+1}$. Hence if $t_{s+1}^{\prime}>0$, then we can use a greedy algorithm to find a $K_{r-s+1}$-tiling $T$ of size $t_{s+1}^{\prime}$ in $G[B_{g}^{1^{\prime}}]$ in time $O(n^{r-s+2})$, as $\mu^{2/3}<\mu^{7/12}$.
@@
Let $K=K^{1}\cup K^{2}$ and $(A_{1}^{2},\ldots,A_{s}^{2},B^{2})=(A_{1}^{1},\ldots,A_{s}^{1},B^{1})-K$. Then $A_{1}^{2},\ldots,A_{s}^{2},B^{2}$ is a balanced partition with $|A_{1}^{2}|=\cdots=|A_{s}^{2}|=\frac{|B^{2}|}{r-s}=n/r-|K|\geq n/r-r^{2}\mu^{2/3}n$, as $t_{i}^{\prime}<8\mu rn+4c$ for each $i\in[s]$, $t_{s+1}^{\prime}<s(r-s+1)\mu^{2/3}n$ and $c=o(n^{1/r^{r+3}})<\mu n$.
The “in particular” part of the lemma can be easily seen from the proof. Indeed, $|B^{1}\setminus V(K)|-(r-s)(n/r-|K^{1}|)=t_{s+1}^{\prime}>0$, and so $K$ contains a copy of $K_{r}$ that has exactly $r-s+1$ vertices of $B^{1}$. ∎
We finish this subsection by proving Lemma 6.14.
@@
    Since $\delta(G[B^{1}])\geq(r-s-1)n/r+t_{s+1}-c\geq(r-s-1)n/r-\mu^{1/3}n$ and $|B_{b}^{1}|\leq 3s\mu^{3/5}n$ (Fact 6.8), any $r-s-1$ vertices of $B^{1}$ have at least $\frac{n}{2r}$ common neighbors in $B^{1}_{g}$.
@@
We start with applying Lemma 6.12 to $G$ with the partition $A_{1}^{1},\ldots,A_{s}^{1},B^{1}$ and the set $I^{*}\subseteq[s+1]$. If the lemma outputs the partition $A_{1}^{1},\ldots,A_{s}^{1},B^{1}$ certifying that no $I^{*}$-admissible $K_{r}$-tilings exist, by Lemma 1.7, $G$ has no $K_{r}$-factor. So it remains to consider the case Lemma 6.12 outputs an $I^{*}$-admissible $K_{r}$-tiling $K^{1}$.
We first apply Lemma 6.13 to find a $K_{r}$-tiling $K\supseteq K^{1}$ of size at most $r^{2}\mu^{2/3}n$ in time $O(n^{r})$ such that $K^{1}\subseteq K$ and $V(G)\setminus V(K)$ has a partition $A_{1}^{2},\ldots,A_{s}^{2},B^{2}$ with $|A_{1}^{2}|=\cdots=|A_{s}^{2}|=\frac{|B^{2}|}{r-s}\geq n/r-r^{2}\mu^{2/3}n$, $A_{i}^{2}\subseteq A_{i}^{1}$ for each $i\in[s]$ and $B^{2}\subseteq B^{1}$. In particular, if $|B^{1}\setminus V(K)|-(r-s)(n/r-|K^{1}|)>0$, then $K$ contains a copy of $K_{r}$ that has exactly $r-s+1$ vertices of $B^{1}$. Then we use Lemma 6.14 repeatedly to find a $K_{r}$-tiling $K^{b}$ whose members are all balanced copies of $K_{r}$ and contain all bad vertices of $B^{1}$ not covered by $K$.
@@
The rest of the proofs split into several cases depending on whether $r-s=2$, $I=[s+1]$, $B_{b}^{1}=\emptyset$ and whether $B^{1}$ is disconnected with two odd components. These simple properties can be checked easily though we need a simple Breath-First Search for the last.
@@
Since $(r-s)n/r-3r^{2}\mu^{3/5}n\leq|B^{3}|\leq(r-s)n/r$, we have $\delta(G[B^{3}])\geq(r-s-1)n/r-3r^{2}\mu^{3/5}n-c\geq\left(1-\frac{1}{r-s}-6r^{3}\mu^{3/5}-\frac{2rc}{n}\right)|B^{3}|$. Moreover, since $B^{1}$ has no $\frac{\lambda}{2}$-sparse $\frac{n}{r}$-set of $G$, it follows that $B^{3}$ has no $\frac{\lambda}{4}$-sparse $\frac{|B^{3}|}{r-s}$-set. Since $r-s\neq 2$, $\mu\ll\lambda$ and $c=o(n^{1/r^{r+3}})$, we can apply Theorem 2.2 to $G[B^{3}]$ and obtain a $K_{r-s}$-factor of $G[B^{3}]$ in time $O(n^{4r^{2}})$. Recall properties (i) and (ii). By applying Lemma 6.3 with $A_{i}^{\prime}=A_{i}^{3}$ for any $i\in[s]$ and $B^{\prime}=B^{3}$, we can find a $K_{r}$-factor of $G[A_{1}^{3}\cup\dots\cup A_{s}^{3}\cup B^{3}]$ in time $O(sn^{4})$. Together with $K$ and $K^{b}$, this gives a $K_{r}$-factor of $G$. ∎
Now suppose $r-s=2$. We use the Edmonds algorithm to check if $B^{3}$ has a perfect matching in time $O(n^{4})$. First suppose it does, and the Edmonds algorithm outputs it. Hence, by applying Lemma 6.3 with $A_{i}^{\prime}=A_{i}^{3}$ for any $i\in[s]$ and $B^{\prime}=B^{3}$, we find a $K_{r}$-factor of $G[A_{1}^{3}\cup\dots\cup A_{s}^{3}\cup B^{3}]$ in time $O(sn^{4})$. Together with $K$ and $K^{b}$, we obtain a $K_{r}$-factor of $G$.
@@
If $r-s=2$ and $I\neq[s+1]$, then we can find a $K_{r}$-factor of $G$ in time $O(n^{r+1})$.
@@
Since $I\neq[s+1]$, we have $|A^{i}\setminus V(K)|-(n/r-|K^{1}|)>0$ for some $i\in[s]$ or $|B^{1}\setminus V(K)|-(r-s)(n/r-|K^{1}|)>0$. Suppose the former one holds. Then there is a large $t_{i}$ for some $i\in[s]\setminus I^{*}$, namely, $t_{i}>c/0.99$. Without loss of generality, $i=1$, and $t_{1}>c/0.99$. By Lemma 6.15 there is an edge in $A_{1}^{3}$, which can be found in $O(n^{2})$. Then extend it to a copy $H_{1}$ of $K_{r}$ with index vector $(2,1,\ldots,1,1)$ by using one vertex from $B^{*}$. We next find a copy of $K_{3}$ in $B^{**}$. Then extend it to a copy $H_{2}$ of $K_{r}$ with index vector $(0,1,\ldots,1,3)$ and vertex-disjoint from $H_{1}$. The running time of the above process is $O(n^{r-2})$.

### v3

@@ 6.5. Auxiliary lemmas
There is an algorithm with running time $n^{O(1)}2^{O(c^{{r}^{r+2}})}$ for $r\geq 4$ (resp. $n^{O(1)}2^{O(c)}$ for $r=3$) which either outputs an $I^{*}$-admissible $K_{r}$-tiling $K^{1}$ or returns $(A_{1}^{1},A_{2}^{1},\ldots,A_{s}^{1},B^{1})$ as a certificate showing that $G$ has no $I^{*}$-admissible $K_{r}$-tiling (and thus also no $K_{r}$-factor).
@@
Suppose that $I^{*}\neq[s+1]$ and Lemma 6.12 outputs an $I^{*}$-admissible $K^{1}$. Then we can always find an $[s+1]$-admissible $K_{r}$-tiling $K$ of size at most $r^{2}\mu^{2/3}n$ in time $O(n^{r})$ such that $K^{1}\subseteq K$ and $V(G)\setminus V(K)$ has a partition $A_{1}^{2},\ldots,A_{s}^{2},B^{2}$ with $|A_{1}^{2}|=\cdots=|A_{s}^{2}|=\frac{|B^{2}|}{r-s}\geq n/r-r^{2}\mu^{2/3}n$, $A_{i}^{2}\subseteq A_{i}^{1}$ for each $i\in[s]$ and $B^{2}\subseteq B^{1}$. In particular, if $|B^{1}\setminus V(K^{1})|-(r-s)(n/r-|K^{1}|)>0$, then $K$ contains a copy of $K_{r}$ that has exactly rs+1r-s+1 vertices of $B^{1}$.
@@
We search for $K_{r}$-tilings $F$ of size at most $g(t_{I^{*}})$ in $G$ such that $\textbf{s}(F)|_{i}=t_{i}$ for each $i\in I^{*}$. By (I3),
$g(t_{I^{*}})=g(t_{I^{*}})\leq(3r(rc/0.33)^{r^{r+1}})^{r}\leq((r^{2}c/0.11)^{r^{r+1}})^{r}=(r^{2}c/0.11)^{r^{r+2}}$
for $r\geq 4$, and
$g(t_{I^{*}})=8t_{I^{*}}\leq(8r)^{r+1}c/0.99$
for $r=3$. (As mentioned at the beginning of Section 6, we only consider $c=o(n^{1/{r^{r+3}}})$ for $r\geq 4$ and $c=o(n)$ for $r=3$. Hence we have $g(t_{I^{*}})=o(n)$ for $r\geq 3$.) By Lemma 1.8, if there is no such $K_{r}$-tiling, then there is no $K_{r}$-factor in $G$. To find such a $K_{r}$-tiling or a certificate showing that none exists, we do the following two steps:
  • Step 1. Search $K_{r}$-tilings of size at most $g(t_{I^{*}})$ in $G$.
@@
Step 1. Recall from Definition 5.1 that if we consider the partition $(A_{1}^{1},\dots,A_{s}^{1},B^{1})$, then there are $y:={r+s\choose s}$ types of copies of $K_{r}$ in $G$. We fix an arbitrary ordering of these $y$ types. Let $x_{i}$ be the number of copies of $i$th-type $K_{r}$ in a $K_{r}$-tiling of size at most $g(t_{I^{*}})$. Thus if we only concern about the types of $K_{r}$’s (that is, identify copies of $K_{r}$ of same type), then the number of distinct $K_{r}$-tilings of size at most $g(t_{I^{*}})$ is equal to the number of solutions of $\sum_{i=1}^{y}x_{i}\leq g(t_{I^{*}})$, i.e. the number of solutions of $\sum_{i=1}^{y}x_{i}+x_{y+1}=g(t_{I^{*}})$ with $x_{1},\dots,x_{y+1}\in\{0\}\cup\mathbb{N}$, which is ${g(t_{I^{*}})+y\choose y}\leq(g(t_{I^{*}})+y)^{y}=O(n^{y})$. For each solution $(x_{1},x_{2},\ldots,x_{y})$ of $\sum_{i=1}^{y}x_{i}\leq g(t_{I^{*}})$, we can decide whether $G$ has a $K_{r}$-tiling of size $C:=\sum_{i=1}^{y}x_{i}$ in time $O(n^{r+1}2^{O(rC)})=O(n^{r+1}2^{O(g(t_{I^{*}}))})$ by Lemma 5.4. Hence the running time of this step is $O(n^{y})\cdot O(n^{r+1}2^{O(g(t_{I^{*}}))})=n^{O(1)}2^{O(g(t_{I^{*}}))}$.
Step 2. Note that checking the slack of a $K_{r}$-tiling takes $O(n)$ time. Since Step 1 returns at most ${g(t_{I^{*}})+y\choose y}=O(n^{y})$ $K_{r}$-tilings, the running time of this step is $O(n^{y})\cdot O(n)=n^{O(1)}$. If there is a $K_{r}$-tiling $F$ obtained in Step 1 satisfying $\textbf{s}(F)|_{i}=t_{i}$ for each $i\in I^{*}$, then we take it as $K^{1}$. Otherwise, by Lemma 1.8, there is no $K_{r}$-factor in $G$, then we halt and output the partition $(A_{1}^{1},\ldots,A_{s}^{1},B^{1})$ of $G$ as a certificate.
Note that the total running time of the above two steps is $n^{O(1)}2^{O(g(t_{I^{*}}))}$, which is $n^{O(1)}2^{O(c^{r^{r+2}})}$ for $r\geq 4$ and $n^{O(1)}2^{O(c)}$ for $r=3$. ∎
@@
Proof.
It suffices to prove that there is an edge even after deleting any set $U\subseteq A_{i}^{1}$ of size $2(xt_{i}-1)$. Since $|A_{i}^{1}|=\frac{n}{r}+t_{i}$, we have $\delta(G[A_{i}^{1}])\geq t_{i}-c$. On the one hand, $e(G[A_{i}^{1}])\geq\frac{1}{2}|A_{i}^{1}|\delta(G[A_{i}^{1}])\geq\frac{1}{2}(\frac{n}{r}+t_{i})(t_{i}-c)\geq\frac{n}{2r}(t_{i}-c)$. On the other hand, $2(xt_{i}-1)\Delta(G[A_{i}^{1}])\leq 2(xt_{i}-1)(3\mu^{1/3}n+2c)\leq 2xt_{i}\cdot 6\mu^{1/3}n$, as $c=o(n)$. Hence there is an edge in $G[A_{i}^{1}\setminus U]$ as long as $\frac{n}{2r}(t_{i}-c)>2xt_{i}\cdot 6\mu^{1/3}n$, that is, $x<1/(2400r\mu^{1/3})$. ∎
@@
Let ${A_{i}^{1}}^{\prime}=A_{i}^{1}\setminus V(K^{1})$ and ${B^{1}}^{\prime}=B^{1}\setminus V(K^{1})$. Let $(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{s+1}^{\prime})$ be the slack of partition $({A_{1}^{1}}^{\prime},\ldots,{A_{s}^{1}}^{\prime},{B^{1}}^{\prime})$. Then we have $t_{i}^{\prime}=0$ for each $i\in I^{*}$ and $\sum_{i\in[s+1]\setminus I^{*}}t_{i}^{\prime}=0$. Note that $|K^{1}|\leq g(t_{I^{*}})$.
For each $i\in[s]\setminus I^{*}$, we have $t_{i}^{\prime}\leq t_{i}+|K^{1}|\leq t_{i}+g(t_{I^{*}})$ (a copy of $K_{r}$ in $K^{1}$ can only make the slack of $A_{i}^{1}$ increase by at most 1). Recall that $|t_{i}|\geq\max\{g(t_{I^{*}}),c/0.99\}$ by (I2). Hence, if $t_{i}\geq\max\{g(t_{I^{*}}),c/0.99\}$, then $t_{i}^{\prime}\leq 2t_{i}$; if $t_{i}\leq\min\{-g(t_{I^{*}}),-c/0.99\}$, then $t_{i}^{\prime}\leq 0$. By Lemma 6.15, if $t_{i}\geq\max\{g(t_{I^{*}}),c/0.99\}$, then $A_{i}^{1}$ has a matching of size at least $(r+2)t_{i}$ and so $A_{i}^{1^{\prime}}$ has a matching of size at least $2t_{i}$. Hence if $t_{i}^{\prime}>0$, then we can find a matching $M_{i}$ of size $t_{i}^{\prime}$ in time $O(n^{4})$ by the Edmonds algorithm.
If $s+1\notin I^{*}$, then we have $t_{s+1}^{\prime}\leq t_{s+1}+(r-s)|K^{1}|$. Hence, if $t_{s+1}\geq g(t_{I^{*}})$, then $t_{s+1}^{\prime}\leq(r-s+1)t_{s+1}<s(r-s+1)\mu^{2/3}n$, as $t_{s+1}<s\mu^{2/3}n$; if $t_{s+1}\leq-g(t_{I^{*}})$, then $t_{s+1}^{\prime}\leq(r-s-1)g(t_{I^{*}})=o(n)$, as $g(t_{I^{*}})=o(n)$. By Lemma 6.9, there are at least $\mu^{7/12}n^{r-s+1}$ copies of $K_{r-s+1}$ in $G[B_{g}^{1}]$, and so $G[B_{g}^{1^{\prime}}]=G[B_{g}^{1}\setminus V(K^{1})]$ contains at least $\mu^{7/12}n^{r-s+1}-r|K^{1}|n^{r-s}\geq(1/2)\mu^{7/12}n^{r-s+1}$ copies of $K_{r-s+1}$. Hence if $t_{s+1}^{\prime}>0$, then we can use a greedy algorithm to find a $K_{r-s+1}$-tiling $T$ of size $t_{s+1}^{\prime}$ in $G[B_{g}^{1^{\prime}}]$ in time $O(n^{r-s+2})$, as $\mu^{2/3}<\mu^{7/12}$.
@@
Let $K=K^{1}\cup K^{2}$ and $(A_{1}^{2},\ldots,A_{s}^{2},B^{2})=(A_{1}^{1},\ldots,A_{s}^{1},B^{1})-K$. Then $A_{1}^{2},\ldots,A_{s}^{2},B^{2}$ is a balanced partition with $|A_{1}^{2}|=\cdots=|A_{s}^{2}|=\frac{|B^{2}|}{r-s}=n/r-|K|\geq n/r-r^{2}\mu^{2/3}n$, as $t_{i}^{\prime}<8\mu rn+4c$ for each $i\in[s]$, $t_{s+1}^{\prime}<s(r-s+1)\mu^{2/3}n$ and $c=o(n)$.
The “in particular” part of the lemma can be easily seen from the proof. Indeed, $|B^{1}\setminus V(K^{1})|-(r-s)(n/r-|K^{1}|)=t_{s+1}^{\prime}>0$, and so $K$ contains a copy of $K_{r}$ that has exactly rs+1r-s+1 vertices of $B^{1}$. ∎
We finish this Section 6.5 by proving Lemma 6.14.
@@
    Since $\delta(G[B^{1}])\geq(r-s-1)n/r+t_{s+1}-c\geq(r-s-1)n/r-\mu^{1/3}n$ and $|B_{b}^{1}|\leq 3s\mu^{3/5}n$ (Fact 6.8), any rs1r-s-1 vertices of $B^{1}$ have at least $\frac{n}{2r}$ common neighbors in $B^{1}_{g}$.
@@
We start with applying Lemma 6.12 to $G$ with the partition $A_{1}^{1},\ldots,A_{s}^{1},B^{1}$ and the set $I^{*}\subseteq[s+1]$. If the lemma outputs the partition $A_{1}^{1},\ldots,A_{s}^{1},B^{1}$ showing that no $I^{*}$-admissible $K_{r}$-tilings exist, by Lemma 1.8, $G$ has no $K_{r}$-factor. So it remains to consider the case Lemma 6.12 outputs an $I^{*}$-admissible $K_{r}$-tiling $K^{1}$.
We first apply Lemma 6.13 to find a $K_{r}$-tiling $K\supseteq K^{1}$ of size at most $r^{2}\mu^{2/3}n$ in time $O(n^{r})$ such that $K^{1}\subseteq K$ and $V(G)\setminus V(K)$ has a partition $A_{1}^{2},\ldots,A_{s}^{2},B^{2}$ with $|A_{1}^{2}|=\cdots=|A_{s}^{2}|=\frac{|B^{2}|}{r-s}\geq n/r-r^{2}\mu^{2/3}n$, $A_{i}^{2}\subseteq A_{i}^{1}$ for each $i\in[s]$ and $B^{2}\subseteq B^{1}$. In particular, if $|B^{1}\setminus V(K^{1})|-(r-s)(n/r-|K^{1}|)>0$, then $K$ contains a copy of $K_{r}$ that has exactly rs+1r-s+1 vertices of $B^{1}$. Then we use Lemma 6.14 repeatedly to find a $K_{r}$-tiling $K^{b}$ whose members are all balanced copies of $K_{r}$ and contain all bad vertices of $B^{1}$ not covered by $K$.
@@
The rest of the proofs is split into several cases depending on whether rs=2r-s=2, $I^{*}=[s+1]$, $B_{b}^{1}=\emptyset$ and whether $B^{1}$ is disconnected with two odd components. These simple properties can be checked easily though we need a simple Breath-First Search for the last.
@@
Since $(r-s)n/r-3r^{2}\mu^{3/5}n\leq|B^{3}|\leq(r-s)n/r$, we have $\delta(G[B^{3}])\geq(r-s-1)n/r-3r^{2}\mu^{3/5}n-c\geq\left(1-\frac{1}{r-s}-6r^{3}\mu^{3/5}-\frac{2rc}{n}\right)|B^{3}|$. Moreover, since $B^{1}$ has no $\frac{\lambda}{2}$-sparse $\frac{n}{r}$-set of $G$, it follows that $B^{3}$ has no $\frac{\lambda}{4}$-sparse $\frac{|B^{3}|}{r-s}$-set. Since $r-s\neq 2$, $\mu\ll\lambda$ and $c=o(n)$, we can apply Theorem 2.2 to $G[B^{3}]$ and obtain a $K_{r-s}$-factor of $G[B^{3}]$ in time $O(n^{4r^{2}})$. Recall properties (i) and (ii). By applying Lemma 6.3 with $A_{i}^{\prime}=A_{i}^{3}$ for any $i\in[s]$ and $B^{\prime}=B^{3}$, we can find a $K_{r}$-factor of $G[A_{1}^{3}\cup\dots\cup A_{s}^{3}\cup B^{3}]$ in time $O(sn^{4})$. Together with $K$ and $K^{b}$, this gives a $K_{r}$-factor of $G$. ∎
Now suppose rs=2r-s=2. We use the Edmonds algorithm to check if $B^{3}$ has a perfect matching in time $O(n^{4})$. First suppose it does, and the Edmonds algorithm outputs it. Hence, by applying Lemma 6.3 with $A_{i}^{\prime}=A_{i}^{3}$ for any $i\in[s]$ and $B^{\prime}=B^{3}$, we find a $K_{r}$-factor of $G[A_{1}^{3}\cup\dots\cup A_{s}^{3}\cup B^{3}]$ in time $O(sn^{4})$. Together with $K$ and $K^{b}$, we obtain a $K_{r}$-factor of $G$.
@@
If rs=2r-s=2 and $I^{*}\neq[s+1]$, then we can find a $K_{r}$-factor of $G$ in time $O(n^{r+1})$.
@@
Since $I^{*}\neq[s+1]$, we have $|A_{i}^{1}\setminus V(K^{1})|-(n/r-|K^{1}|)>0$ for some $i\in[s]$ or $|B^{1}\setminus V(K^{1})|-(r-s)(n/r-|K^{1}|)>0$. Suppose the former one holds. Then there is a large $t_{i}$ for some $i\in[s]\setminus I^{*}$, namely, $t_{i}>c/0.99$. Without loss of generality, set $i=1$, and then $t_{1}>c/0.99$. By Lemma 6.15 there is an edge in $A_{1}^{3}$, which can be found in $O(n^{2})$. Then we extend it to a copy $H_{1}$ of $K_{r}$ with index vector $(2,1,\ldots,1,1)$ by using one vertex from $B^{*}$. We next find a copy of $K_{3}$ in $B^{**}$, and then extend it to a copy $H_{2}$ of $K_{r}$ with index vector $(0,1,\ldots,1,3)$ and it is vertex-disjoint from $H_{1}$. The running time of the above process is $O(n^{r-2})$.